he following is summary of information presented on the financial statements of a company on December 31, 2015. Account 2015 2014 Net Sales Revenue $600,000 $500,000 Cost of Goods Sold 450,000 400,000 Gross Profit $150,000 $100,000 Selling Expenses 50,000 50,000 Net income before income tax expense $100,000 $50,000 Income tax expense 35,000 18,000 Net Income $65,000 $32,000

What would a horizontal analysis report show with respect to net income?

both net income before income tax expense and net income are 45.45% of net sales revenue

a $25,000 increase in net income

a 45.45% increase in net income

a 103% increase in net income
Answer #1

Changes in net income :

2014 net income = 32000

2015 net income = 65000

Increase in net income = 65000-32000 = 33000

Changes in percentage = 33000*100/32000 = 103%

so answer is d) a 103% increase in net income